find the derivative of abs(x-3) when  x = -1 with correct steps

Grade:12

2 Answers

Sourabh Singh IIT Patna
7 years ago
Hii
you have to draw the curve and once you draw it you will see that the slope of the tangent is -1
Ajay
7 years ago
By definition the modulus function is defined as 
\left | x-3 \right | = -(x-3)\ if\ x< 3 \\
= -x+3 
Hence derivative  = -1 for all values of x less than 3 which includes -1